How many gallons of water that is 74% alcohol must be mixed with 5 gallons of one that is 90% alcohol in order to obtain a mixture that is 84% alcohol?
Answer by richwmiller(17219)

We have 5 gal of 90% solution
We want to add some 74% solution to make 84% solution
0.74*x+0.9*5=0.84(5+x)
0.74*x+4.5=4.2+0.84x
0.74x-0.84x=4.2-4.5
-0.1x=-0.3
x=3 gal at 74% (2.22) is added to 5 gal at 90%(4.5) to earn 84% of 8 gal(6.72)
check
0.74*3+0.9*5=0.84(5+3)
2.22+4.5=0.84(8)
6.72=6.72
ok